OpenStudy (anonymous):

Solve linear congruence \[6x≡9 mod 15\]

OpenStudy (anonymous):

@jim_thompson5910 @Hero

jimthompson5910 (jim_thompson5910):

One way to do it: x = 1, 6x = 6*1 = 6 (mod 15) x = 2, 6x = 6*2 = 12 (mod 15) x = 3, 6x = 6*3 = 18 = 3 (mod 15) x = 4, 6x = 6*4 = 24 = 9 (mod 15) So we can stop here

OpenStudy (anonymous):

the method i see in my book is to find the inverse congruence class, which i am very confused on can you please explain the inverse congruence method? thanks

jimthompson5910 (jim_thompson5910):

To find the inverse of 6 mod 15, you would solve the congruence below 6y = 1 (mod 15) for y

jimthompson5910 (jim_thompson5910):

but the problem is that gcd(6,15) = 3 and this doesn't divide 1 So 6 doesn't have an inverse mod 15 ie, there are no solutions to 6y = 1 (mod 15)

jimthompson5910 (jim_thompson5910):

so you'll be spending quite a bit of time looking for the inverse, but you'll never find it (because it doesn't exist in this case)

OpenStudy (anonymous):

the first step in the book was to divide the whole equation by 3, which would give \[2x≡3mod5\]and then it somehow jumps to\[x≡4mod5\]which confused the bajeezus out of me

jimthompson5910 (jim_thompson5910):

oh it's using the rule that if If ma = mb (mod n), and if d=GCD(m,n), then a = b (mod n/d)

OpenStudy (anonymous):

ok i get that so far.

jimthompson5910 (jim_thompson5910):

6x = 9 (mod 15) 3*2x = 3*3 (mod 3*5) 2x = 3 (mod 5)

jimthompson5910 (jim_thompson5910):

now you can find the inverse because you can solve the following for y 2y = 1 (mod 5) the solution is y = 3 since 2*3 = 6 which is 1 higher than 5 (so 6 = 1 (mod 5)) so... 2*3 = 1 (mod 5) 2*3*3 = 1*3 (mod 5) 2*9 = 3 (mod 5) 2*4 = 3 (mod 5) So the solution to 2x = 3 (mod 5) is x = 4 (mod 5)

OpenStudy (anonymous):

ok that makes a lot of sense now. because 9 is actually 5+4, so 9 mod 5 is the same as 4 mod 5?

jimthompson5910 (jim_thompson5910):

exactly

jimthompson5910 (jim_thompson5910):

or another way to think of it 9/5 = 1 remainder 4 that remainder is exactly what the mod is saying